I don't know if this is the correct topic but here is my question below:
The floor of a hall is covered by square tiles all of the same size laid side by side to form a rectangle of 105 tiles b 84 tiles. If you draw a line from one corner of the hall to the opposite corner of the hall, how many tiles will the line intersect?

Answer by rothauserc(4718)

105 and 84 have a common factor of 21
:
the diagonal will cross 21 times as many tiles as the diagonal of a 105/21 = 5 by 84/21 = 4 tile rectangle.
:
Since 5 and 4 have no common factor, the answer for the small
rectangle is 5 + 4 - 1 = 8.
:
***********************************
The final answer is 8*21 = 168
***********************************
:
Note For the 5 by 4 tile rectangle start at one corner with one tile. What happens every time the diagonal crosses either a horizontal or vertical line? Why do we need 5 and 4 to have no common factor?
: